The length of a rectangular plot of land is 4 times the width. If the perimeter is 2,000 feet, find the dimensions of the plot.
Answer by mananth(16949)

let width be x
length = 4x
2(4x+x)=2000
5x =2000/2
5x= 1000
/5
x= 200 feet width
length = 4x = 800 feet
...
m.ananth@hotmail.ca